OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = eight divided by x squared + 2 please help and explain

OpenStudy (anonymous):

You are looking for two functions f(x) and g(x). With f(x) being the first function and then you incorporate g(x) into that function to come up with y. Example comming

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

\[y = 4/(x+2).....if f(x) =4/x ....then... g(x)=(x+2)\]

OpenStudy (anonymous):

Does this help?

OpenStudy (anonymous):

well it gives me f(x) and g(x) like do I find the f(x) and g(x) in each?

OpenStudy (anonymous):

Actually I gave you the equation for y. The equation should have been written above as f(g(x) = 8 / x^2+2.

OpenStudy (anonymous):

Now solve for f(x) and g(x)

OpenStudy (anonymous):

so f(x) is 8 and g(x) is x^2+2 ??

OpenStudy (anonymous):

Close...f(x) = 8/x and g(x) is x^2+2. Then f(g(x)) = 8/x^2+2. Because if f(g(x)) we take the equation for f(x) and plug in g(x) for x. Make sense?

OpenStudy (anonymous):

thanks how about this one this is the last I tried doing it but square root threw me off Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x2 - 3 and g(x) = square root of quantity three plus x

OpenStudy (anonymous):

Do you know how do find the inverse of a function?

OpenStudy (anonymous):

well would it be x/3 + square 3+x? for f(x(g))

OpenStudy (anonymous):

Never mind a misread the question. Take f(x) = x^2 -3 and plug in g(x) for the x value what do you get?

OpenStudy (anonymous):

explain?

OpenStudy (anonymous):